A object whose mass is 0.0050 kg is in simple harmonic motion with a period of 0.040 sec and

an amplitude of 10.0 mm. a) What is it’s maximum acceleration? b) What is the maximum force

on the object? c) What is it’s acceleration when it is 5.00 mm from it’s equilibrium position?

d) What is the objects total energy?


1
Expert's answer
2021-05-25T10:10:18-0400

a) "a=\frac{4\pi^2(0.01)}{0.04^2}=247\frac{m}{s^2}"

b) "F=0.005(247)=1.23\ N"

c) "a'=157^2(0.005)=123 \frac{m}{s^2}"

d) "E=0.5(0.005)(157)^2(0,01)^2=0.0062\ J"